Find the value of a such that 2 is a root of the polynomial 2.4 – 23 + ax - 1
Answer #1

Let f(x) = 2x⁴ - x³ + ax - 1

If 2 is a root of the polynomial f(x), then f(2) = 0

f(2) = 0

2(2)⁴ - (2)³ + a(2) - 1 = 0

2(16) - 8 + 2a - 1 = 0

32 - 9 + 2a = 0

23 + 2a = 0

2a = -23

a = (-23/2)

Hence the required value of a is (-23/2).